OpenStudy (anonymous):

if f(x)=2x-3 and g(x)=2x^2+1 find A.f(g(2)) B. g(f(x)) C. does f(x) have and inverse if so find the inverse of f(x)

OpenStudy (anonymous):

a. g(2) = 2*4+1 = 9 f(g(2)) = 2*9 - 3 = 15

OpenStudy (anonymous):

oh ok thank you

OpenStudy (anonymous):

b. g(f(x)) = 2 (2x-3)^2 +1 = 2 (4x^2 -12x +9)+1 = 8x^2 - 24x +19

OpenStudy (anonymous):

im a lil bit confused about a

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

so when you found g(2) you pluugged it infor f(x)?

OpenStudy (anonymous):

yes, because g(2) = x for f(x)

OpenStudy (anonymous):

oh ok thanks